OpenStudy (anonymous):

An equation is shown below: 9x - y = -2 Part C: Write an equation that can be paired with the given equation in order to form a system of equations that is inconsistent. (2 points)

OpenStudy (jdoe0001):

hmmm when is a system of equations "inconsistent"?

OpenStudy (anonymous):

When it's not linear?

OpenStudy (anonymous):

9x-y=-2 -9x -9x y=-9x-2 this is what i got so far

OpenStudy (jdoe0001):

well... both are likley linear equations is inconsistent if it has no solutions not solutions for a system of equations with linear equations means the lines never touch each other and that only happens when both lines are parallel

OpenStudy (anonymous):

So how do I make an equation right there that is inconsistent? Make it quick please I'm going to my dad's birthday dinner soon.

OpenStudy (jdoe0001):



OpenStudy (jdoe0001):

do you know how to find slopes in a line?

OpenStudy (anonymous):

Rise/Over run... Count up x and count right y.,

OpenStudy (jdoe0001):

well.. yes... if you just solve the one you have there, for say "y" what would you end up?

OpenStudy (jdoe0001):

hmm anyhow.. if both lines have the same "slope" that means they're both parallel so if you solve that one for "y" and put it in mx+b slope-intercept form.... keep the slope, change the y-intercept and you'd end up with a parallel line to it

OpenStudy (jdoe0001):



OpenStudy (anonymous):

So what would the equation be that would make it inconsistent?

OpenStudy (jdoe0001):

well.. what did you get when solving for "y" at 9x - y = -2 ?

OpenStudy (anonymous):

Well, 9(2) - y = -2 would just be 20.

OpenStudy (anonymous):

y=20.

OpenStudy (jdoe0001):

hmmm? well.. 20? what happened to the "x" though?